### v2

# The multiplicative Kowalski-Słodkowski Theorem for Hermitian algebras

Abstract.
We prove, for Hermitian algebras, the multiplicative version of the Kowalski-Słodkowski Theorem which identifies the characters among the collection of all complex valued functions on a Banach algebra $A$ in terms of a spectral condition. Specifically, we show that, if $A$ is a Hermitian algebra, and if $\phi:A\mapsto\mathbb{C}$ is a continuous function satisfying $\phi(x)\phi(y)\in\sigma(xy)$ for all $x,y\in A$ (where $\sigma$ denotes the spectrum), then either $\phi$ or $\phi$ is a character of $A$; of course the converse holds as well. Our proof depends fundamentally on the existence of positive elements and square roots in these algebras.
Key words and phrases:
Hermitian algebra, Banach algebra, spectrum, spectral radius, linear function, multiplicative function, character
2020 Mathematics Subject Classification:
46H05, 46H15, 47A10
1 Introduction
Let $A$ be a complex Banach algebra with the identity and zero denoted by $\mathbf{1}$ and $\mathbf{0}$ respectively. Let $G(A)$ be the invertible group of $A$, and $G_{\mathbf{1}}(A)$ be the connected component of $G(A)$ containing $\mathbf{1}$. For $x\in A$, denote by $\sigma(x):=\left\{\lambda\in\mathbb{C}:\lambda\mathbf{1}-x\not\in G(A)\right\}$ the spectrum of $x$. A nonzero function $\phi:A\to\mathbb{C}$ (not assumed to be linear) is said to be multiplicative if $\phi(xy)=\phi(x)\phi(y)$ for all $x,y\in A$. If $\phi:A\to\mathbb{C}$ (not assumed to be linear or multiplicative) satisfies $\phi(x)\phi(y)\in\sigma(xy)$ for all $x,y\in A,$ then $\phi$ is called spectrally multiplicative. Of course, any nonzero function $\phi:A\to\mathbb{C}$ which is both linear and multiplicative is called a character of $A$. A classical result, dating back to the late 1960’s identifies the characters of $A$ among the linear functionals on $A$:
Theorem 1.1 (Gleason-Kahane-Żelazko Theorem).
Let $A$ be a complex unital Banach algebra, and suppose $\phi:A\rightarrow\mathbb{C}$ is linear. Then $\phi$ is a character of $A$ if and only if $\phi(x)\in\sigma(x)$ for each $x\in A$.
For examples showing that Theorem 1.1 does not extend to real Banach algebras look at [7, p.43]. However, in the real case, if one replaces the requirement $\phi(x)\in\sigma(x)$ for each $x\in A$ in the statement of Theorem 1.1 by the two conditions $\phi(\mathbf{1})=1$ and $\phi(x)^{2}+\phi(y)^{2}\in\sigma(x^{2}+y^{2})$ for all commuting pairs $x,y\in A$, then $\phi$ is a character of $A$; this is due to Kulkarni [3].
A stronger result than Theorem 1.1, obtained by Kowalski and Słodkowski in [2], identifies the characters among all complex-valued functions on $A$ via a spectral condition:
Theorem 1.2 (Kowalski-Słodkowski Theorem).
Let $A$ be a complex unital Banach algebra. Then a function $\phi:A\rightarrow\mathbb{C}$ is a character of $A$ if and only if $\phi$ satisfies
  • (i)
    $\phi(\mathbf{0})=0$,
  • (ii)
    $\phi(x)-\phi(y)\in\sigma(x-y)$ for every $x,y\in A$.
It is obvious, in Theorem 1.2, that (i) and (ii) together can be replaced by the single requirement $\phi(x)+\phi(y)\in\sigma(x+y)$ for every $x,y\in A.$
In [4], Li, Peralta, Wang and Wang established spherical variants of the Gleason-Kahane-Żelazko and Kowalski-Słodkowski Theorems which they then use to prove that every weak-2-local isometry between two uniform algebras is a linear map. In the following $\mathbb{T}$ denotes the unit circle in $\mathbb{C}$.
Theorem 1.3 (Spherical Gleason-Kahane-Żelazko Theorem; [4, Proposition 2.2]).
Let $A$ be a complex unital Banach algebra, and suppose $\phi:A\rightarrow\mathbb{C}$ is a continuous linear functional. If $\phi(x)\in\mathbb{T}\sigma(x)$ for each $x\in A$, then $\overline{\phi(\mathbf{1})}\phi$ is a character of $A$.
Theorem 1.4 (Spherical Kowalski-Słodkowski Theorem; [4, Proposition 3.2]).
Let $A$ be a complex unital Banach algebra, and suppose $\phi:A\rightarrow\mathbb{C}$ satisfies
  • (i)
    $\phi(\lambda x)=\lambda\phi(x$) for each $x\in A$, $\lambda\in\mathbb{C}$,
  • (ii)
    $\phi(x)-\phi(y)\in\mathbb{T}\sigma(x-y)$ for every $x,y\in A$.
Then $\lambda_{0}\phi$ is a character for some $\lambda_{0}\in\mathbb{T}$.
In [8] Oi shows that if one relaxes the homogeneity condition (i) in Theorem 1.4 to $\phi(\mathbf{0})=0$, then $\phi$ is either complex-linear or conjugate linear and $\overline{\phi(\mathbf{1})}\phi$ is multiplicative.
The Gleason-Kahane-Żelazko and Kowalski-Słodkowski Theorems naturally lead to considering the scenario where additivity is sacrificed for multiplicativity. The first major result pertaining to this question was obtained by Maouche:
Theorem 1.5 (Maouche, [6]).
Let $A$ be a Banach algebra, and let $\phi:A\rightarrow\mathbb{C}$ be a multiplicative function satisfying $\phi(x)\in\sigma(x)$ for each $x\in A$. Then, corresponding to $\phi$, there exists a unique character of $A$ which agrees with $\phi$ on $G_{\mathbf{1}}(A)$.
Maouche concludes with an example (where $A=C[0,1]$) showing that $\phi$ in his theorem might not be a character itself. In [5] (stated here as Theorem 1.7) it is shown that the missing ingredient in his example is continuity of $\phi$. So this prompts the question: If continuity of $\phi$ is added to the hypothesis in Maouche’s Theorem, does it then follow that $\phi$ is a character of $A$? If this is indeed the case, then we suspect that the proof might not be easy. Observe also that continuity is not required in the hypotheses of either Theorem 1.1 or Theorem 1.2, and is part of the conclusion rather than the assumption. Substantial progress on the aforementioned conjecture has been made for algebras with involution. Theorems 1.6 and 1.7 stated below are the latest results in this context.
Theorem 1.6 ([11, Corollary 3.8]).
Let $A$ be a $C^{\star}$-algebra and let $\phi:A\to\mathbb{C}$ be a continuous spectrally multiplicative function. Then either $\phi$ or $\phi$ is a character of $A$.
Theorem 1.7 ([5, Theorem 1.2]).
Let $A$ be a Hermitian algebra and let $\phi:A\to\mathbb{C}$ be a continuous multiplicative function such that $\phi(x)\in\sigma(x)$ for each $x\in A$. Then $\phi$ is a character of $A$.
The aim of the current paper is to simultaneously improve on Theorems 1.6 and 1.7. Specifically, Theorem 1.6 holds for all Hermitian algebras; of course this means that the multiplicativity requirement on $\phi$ in Theorem 1.7 can be relaxed to spectral multiplicativity with the conclusion that either $\phi$ or $\phi$ is a character of $A$ (in Theorem 1.7 the fact that $\phi(\mathbf{1})=1$ is immediate from the hypothesis). Expectedly, our proof combines techniques used in [5] and [11], but we also need the following result, Theorem 1.8, from [10] since there is currently no multiplicative Kowalski-Słodkowski version of Theorem 1.5. Of course, from this emerges another conjecture. If $A$ is a unital Hermitian algebra, we denote by $S_{A}$ the self-adjoint elements of $A$.
Theorem 1.8 ([10, Theorem 3.7, Theorem 3.9]).
Let $A$ be a Hermitian algebra and let $\phi:A\to\mathbb{C}$ be a continuous spectrally multiplicative function satisfying $\phi(\mathbf{1})=1$. Then the formula
$\psi_{\phi}(x):=\phi\left(\operatorname{Re}(x)\right)+i\phi\left(\operatorname{Im}(x)\right)$
defines a character of $A$. Further, $\psi_{\phi}$ agrees with $\phi$ on $G_{\mathbf{1}}(A)\cup S_{A}.$
Remark.
In [10] Theorem 1.8 was proved for $C^{\star}$-algebras. A close inspection of the proof shows that the arguments presented there rely only on the fact that self-adjoint elements have real spectra, and the Lie-Trotter formula which says that for $x,y$ in any unital Banach algebra AA
$\lim_{n\rightarrow\infty}\left(e^{x/n}e^{y/n}\right)^{n}=e^{x+y}.$
Theorem 1.8 is therefore also valid for Hermitian algebras. In [1] Escolano, Peralta, and Villena derive a number of Lie-Trotter formulae for Jordan-Banach algebras and they then use this to derive the multiplicative version of the Gleason-Kahane-Żelazko Theorem for $JB^{\star}$-algebras [1, Theorem 4.3].
2 Preliminaries
Throughout the remainder of this paper $A$ will be a complex unital Hermitian algebra. By definition the positive and strictly positive elements of $A$ are those self-adjoint elements of $A$ whose spectra are contained in $\mathbb{R}^{\geq 0}$ and $\mathbb{R}^{>0}$, respectively. For $a,b\in S_{A}$ we will write $a\leq b$ (resp. $a<b$) if $0\leq b-a$ (resp. 0<ba0<b-a). If $\rho$ denotes the spectral radius, then we define the Pták functional by $s(x)\coloneqq\sqrt{\rho(x^{\star}x)}$, and we recall that $s$ is both subadditive and submultiplicative. Further, since the Pták functional dominates the spectral radius it is, as in the case with the norm, useful for detecting invertibility. From [9, Theorem 11.20], if $0<a$, then $a$ has a positive square root. That is, there exists $y\in A$ such that $y^{2}=a$, $\sigma(y)\subset(0,\infty)$ and $y=y^{\star}$. In particular, this square root is not given by a continuous functional calculus but rather by the holomorphic functional calculus viz
(1) $y=\frac{1}{2\pi i}\int\limits_{\Gamma}e^{\frac{1}{2}\log\lambda}\left(\lambda\mathbf{1}-a\right)^{-1}\,d\lambda,$
where $\log\lambda$ is the principal branch of the complex logarithm, and $\Gamma$ is a smooth, positively oriented, contour surrounding $\sigma(a)$, not separating 0 from infinity. The first two properties of $y$ in (1) are obvious; that $y=y^{\star}$ is more interesting, and a nice exposition of the argument can be found in [9]. We denote this positive square root by $y=\sqrt{a}$.
Hermitian algebras are not necessarily semisimple, and in these cases the involution may not be continuous. To deal with this situation (in the forthcoming section) denote by $\operatorname{rad}(A)$ the radical of $A$, and let $\pi:A\to A/\operatorname{rad}(A)$ be the canonical homomorphism. It is elementary to show that $\pi(x)^{\star}:=\pi(x^{\star})$ is a well-defined involution on $A/\operatorname{rad}(A)$. Since $\sigma(x)=\sigma(\pi(x))$ for all $x\in A$ it follows that $A/\operatorname{rad}(A)$ is a semisimple Hermitian algebra, and necessarily the involution here is continuous. Moreover, the above spectral equality implies that if $x\in A$ and $x$ is normal modulo $\operatorname{rad}(A)$ i.e $x^{\star}x-xx^{\star}\in\operatorname{rad}(A),$ then $s(x)=\rho(x).$
3 The multiplicative Kowalski-Słodkowski Theorem for Hermitian algebras
To simplify our arguments we shall initially assume that $\phi$ satisfies $\phi(\mathbf{1})=1$. We start with a few easy but somewhat non-trivial lemmas:
Lemma 3.1.
Let $\left(a_{n}\right)$ and $\left(b_{n}\right)$ be sequences in $S_{A}$ such that $a_{n}^{2}<b_{n}^{2}$ for all $n\in\mathbb{N}$. If $\left(v_{n}\right)$ is a sequence in $A$ such that $s(b_{n}v_{n})\to 0$, then $s(a_{n}v_{n})\to 0$.
Proof.
It suffices to show that $s(a_{n}v_{n})^{2}\leq s(b_{n}v_{n})^{2}$ for each $n$. Observe that
$s(a_{n}v_{n})^{2}=\rho((a_{n}v_{n})^{\star}a_{n}v_{n})=\rho(v_{n}^{\star}a_{n}^{2}v_{n}).$
Since $0<b_{n}^{2}-a_{n}^{2}$ it follows, using the Shirali-Ford Theorem, that $v_{n}^{\star}a_{n}^{2}v_{n}\leq v_{n}^{\star}b_{n}^{2}v_{n}$ for each $n$. Moreover, $v_{n}^{\star}b_{n}^{2}v_{n}\leq\rho(v_{n}^{\star}b_{n}^{2}v_{n})\mathbf{1}$ and hence $\rho(v_{n}^{\star}a_{n}^{2}v_{n})\leq\rho(v_{n}^{\star}b_{n}^{2}v_{n}),$ for every $n\in\mathbb{N}.$ But then,
$s(a_{n}v_{n})^{2}=\rho(v_{n}^{\star}a_{n}^{2}v_{n})\leq\rho(v_{n}^{\star}b_{n}^{2}v_{n})=s(b_{n}v_{n})^{2}.$
Lemma 3.2.
Let $a\in S_{A}$, and let $(v_{n})$ be a sequence in $A$ such that $s\left(\sqrt{a^{2}+\frac{1}{n^{2}}\mathbf{1}}\,v_{n}\right)\to 0$. Then $s(av_{n})\to 0$.
Proof.
Obviously, from the assumption, $s\left(2\sqrt{a^{2}+\frac{1}{n^{2}}\mathbf{1}}\,v_{n}\right)\to 0$. Observe that for each $n\in\mathbb{N}$
$0<\left(\sqrt{a^{2}+\frac{1}{n^{2}}\mathbf{1}}-a\right)^{2}<\left(2\sqrt{a^{2}+\frac{1}{n^{2}}\mathbf{1}}\right)^{2}.$
Hence, Lemma 3.1 says that $s\left(\left(\sqrt{a^{2}+\frac{1}{n^{2}}\mathbf{1}}-a\right)v_{n}\right)\to 0$. Using the subadditivity of $s$, we get that $s(av_{n})\to 0$. ∎
Since $xe^{x}$ is the limit of a sequence of polynomials in $x$ with real coefficients we have the following easy
Lemma 3.3.
AA be a Hermitian algebra and suppose $x\in S_{A}$. Then $xe^{x}$ is self-adjoint modulo $\operatorname{rad}(A)$.
Lemma 3.3 can now be used to obtain the next
Lemma 3.4.
Let $0\leq a$. Then
  • (a)
    $\lim\limits_{n\to\infty}s\left(\sqrt{a+\frac{1}{n^{2}}\mathbf{1}}\,e^{-n\sqrt{a+\frac{1}{n^{2}}\mathbf{1}}}\right)=0,$
  • (b)
    $\lim\limits_{n\to\infty}s\left(\sqrt{a+\frac{1}{n^{2}}\mathbf{1}}\left(\mathbf{1}+in\sqrt{a+\frac{1}{n^{2}}\mathbf{1}}\right)^{-1}\right)=0.$
Proof.
Let $(a_{n})$ be a sequence in $A$ with $0<a_{n}$ for each $n\in\mathbb{N}$, and suppose further that $\{a_{n}\colon n\in\mathbb{N}\}$ is a commutative subset of $A$. It follows that the bicommutant $B$ of $\{a_{n}\colon n\in\mathbb{N}\}$ is a Hermitian subalgebra of $A$ preserving the spectra of elements. Consequently, the set $\mathfrak{M}$ of characters of $B$ is nonempty. With $b_{n}=a_{n}e^{-na_{n}}$ we have, using Lemma 3.3, that each $b_{n}$ is self-adjoint modulo $\operatorname{rad}(B)$ from which it follows that
$s(b_{n})=\rho(b_{n})=\sup\left\{\chi(a_{n})e^{-n\chi(a_{n})}\colon\chi\in\mathfrak{M}\right\}\leq\sup\left\{te^{-nt}\colon t\geq 0\right\}\leq\frac{1}{en}.$
Therefore, $s(b_{n})\to 0$. Now let $c_{n}=a_{n}(\mathbf{1}+ina_{n})^{-1}$, so that $c_{n}$ is normal and
$s(c_{n})=\rho(c_{n})=\sup\left\{\frac{\chi(a_{n})}{\lvert 1+in\chi(a_{n})\rvert}\colon\chi\in\mathfrak{M}\right\}\leq\sup\left\{\frac{t}{\lvert 1+int\rvert}\colon t\geq 0\right\}\leq\frac{1}{n}.$
Thus, $s(c_{n})\to 0$. To get the result, let $a_{n}=\sqrt{a+\frac{1}{n^{2}}\mathbf{1}}$ for each $n\in\mathbb{N}$. ∎
Lemma 3.5.
Let $\phi$ be a continuous spectrally multiplicative function on $A$ with $\phi(\mathbf{1})=1$. If $\psi_{\phi}(x)=0$, then $\phi(x)=0$.
Proof.
To simplify notation write $R=\operatorname{Re}(x)$ and $I=\operatorname{Im}(x)$. Then $R$ and $I$ are self-adjoint and $\psi_{\phi}(R)=\psi_{\phi}(I)=0$. For each $n\in\mathbb{N}$, let
$W_{n}\coloneqq e^{-n\sqrt{R^{2}+I^{2}+\frac{2}{n^{2}}\mathbf{1}}}.$
Hence,
$\phi(W_{n})=\psi_{\phi}(W_{n})=e^{-\sqrt{2}}.$
Since $\phi$ is spectrally multiplicative it then follows that
$\phi(x)e^{-\sqrt{2}}=\phi(x)\phi(W_{n})\in\sigma(xW_{n}),\text{ for each }n\in\mathbb{N}.$
From Lemma 3.4, we have that $s\left(\sqrt{R^{2}+I^{2}+\frac{2}{n^{2}}\mathbf{1}}\,W_{n}\right)\to 0$. Notice that
${R^{2}+\frac{1}{n^{2}}\mathbf{1}}<{R^{2}+I^{2}+\frac{2}{n^{2}}\mathbf{1}},$
so Lemma 3.1 says that $s\left(\sqrt{R^{2}+\frac{1}{n^{2}}\mathbf{1}}\,W_{n}\right)\to 0$. Now Lemma 3.2 implies that $s(RW_{n})\to 0$. Similarly, $s(IW_{n})\to 0$. Therefore,
$s(xW_{n})=s(RW_{n}+iIW_{n})\leq s(RW_{n})+s(IW_{n})\to 0.$
From earlier
${\lvert\phi(x)\rvert}e^{-\sqrt{2}}\leq\rho(xW_{n})\leq s(xW_{n}),$
which means $\lvert\phi(x)\rvert\leq e^{\sqrt{2}}s(xW_{n})$. Taking limits as $n\to\infty$ yields $\phi(x)=0$. ∎
Lemma 3.6.
Let $\phi$ be a continuous spectrally multiplicative function on $A$ with $\phi(\mathbf{1})=1$. If $\alpha\in\mathbb{C}$, and $x\in A$ satisfies $\psi_{\phi}(x)=0$, then $\phi(\alpha\mathbf{1}+x)=c_{\alpha}\alpha$ for some $c_{\alpha}\in[0,1]$.
Proof.
If $\alpha=0$ then the result follows from Lemma 3.5, so take $\alpha\neq 0$. Using $W_{n}$ as before, let $Y_{n}=\frac{1}{\alpha}xW_{n}$ and set $c_{\alpha}=\phi(\alpha\mathbf{1}+x)/\alpha$. Then,
${c_{\alpha}}{e^{-\sqrt{2}}}=\frac{1}{\alpha}\phi(\alpha\mathbf{1}+x)\phi(W_{n})\in\sigma(W_{n}+Y_{n}).$
Also, $s(Y_{n})\to 0$ from the previous proof. Suppose, for the sake of contradiction, that $c_{\alpha}\not\in[0,1]$ whence $c_{\alpha}e^{-\sqrt{2}}\not\in\left[0,e^{-\sqrt{2}}\right]$. From the definition of $W_{n}$ one observes that $\sigma(W_{n})\subseteq[0,e^{-\sqrt{2}}]$ for each $n\in\mathbb{N}$. Then we must have that
$\mathbf{1}-Y_{n}\left(c_{\alpha}e^{-\sqrt{2}}\mathbf{1}-W_{n}\right)^{-1}\not\in G(A),$
for otherwise $c_{\alpha}e^{-\sqrt{2}}\mathbf{1}-W_{n}-Y_{n}\in G(A)$ – a contradiction. But this then means that
$s\left(Y_{n}\left(c_{\alpha}e^{-\sqrt{2}}\mathbf{1}-W_{n}\right)^{-1}\right)\geq 1\mbox{ for all }n\in\mathbb{N}.$
Furthermore, $\left(c_{\alpha}e^{-\sqrt{2}}\mathbf{1}-W_{n}\right)^{-1}$ is normal modulo $\operatorname{rad}(A)$ for each $n$, and hence
$s\left(\left(c_{\alpha}e^{-\sqrt{2}}\mathbf{1}-W_{n}\right)^{-1}\right)=\rho\left(\left(c_{\alpha}e^{-\sqrt{2}}\mathbf{1}-W_{n}\right)^{-1}\right)\leq\frac{1}{\operatorname{dist}\left(\left[0,e^{-\sqrt{2}}\right],c_{\alpha}e^{-\sqrt{2}}\right)}.$
Since $s$ is submultiplicative we get that $s\left(Y_{n}\left(c_{\alpha}e^{-\sqrt{2}}\mathbf{1}-W_{n}\right)^{-1}\right)\to 0$, which contradicts the fact that it must be at least $1$ for all $n$. Thus, $\phi(\alpha\mathbf{1}+x)=c_{\alpha}\alpha$, with $c_{\alpha}\in[0,1]$. ∎
Lemma 3.7.
Let $\phi$ be a continuous spectrally multiplicative function $A$ with $\phi(\mathbf{1})=1$. If $\alpha\in\mathbb{C},$ and $x\in A$ satisfies $\psi_{\phi}(x)=0$, then $\phi(\alpha\mathbf{1}+x)\in\{0,\alpha\}.$
Proof.
For each $n\in\mathbb{N}$ let $V_{n}:=\left(\mathbf{1}+in\sqrt{R^{2}+I^{2}+\frac{1}{n^{2}}\mathbf{1}}\right)^{-1}.$ Again using Lemma 3.1, Lemma 3.2 and Lemma  3.4, we have that
$\lim_{n}s\left(\sqrt{R^{2}+I^{2}+\frac{1}{n^{2}}\mathbf{1}}\,V_{n}\right)={0}\implies\lim_{n}s\left(RV_{n}\right)=0.$
Similarly $\lim_{n}s\left(IV_{n}\right)=0$. Observe that each $V_{n}$ belongs to $G_{\mathbf{1}}(A)$, whence it follows that $\phi(V_{n})=\psi_{\phi}(V_{n})=1/2-i/2$. Let $\alpha\neq 0$. From Lemma 3.6, we have that $\phi(\alpha\mathbf{1}+x)=c_{\alpha}\alpha$, with $c_{\alpha}\in[0,1]$. To obtain the result we have to show that $c_{\alpha}\in\{0,1\}$: For the sake of a contradiction assume that $0<c_{\alpha}<1$. If we set $Z_{n}:=\frac{1}{\alpha}xV_{n}=\frac{1}{\alpha}(R+iI)V_{n}$, then
(2) $c_{\alpha}(1/2-i/2)=\frac{1}{\alpha}\phi(\alpha\mathbf{1}+x)\phi(V_{n})\in\sigma\left(V_{n}+Z_{n}\right).$
The first paragraph of the proof shows that $\lim_{n}s\left(Z_{n}\right)=0$, and (2) shows that $c_{\alpha}(1/2-i/2)\mathbf{1}-V_{n}-Z_{n}\notin G(A)$. From the definition of $V_{n}$ together with the spectral mapping theorem, and the fact that $0<\sqrt{R^{2}+I^{2}+\frac{1}{n^{2}}\mathbf{1}}$ we have that $\sigma\left(V_{n}\right)\subseteq C_{r}$, where $C_{r}$ is the circle in $\mathbb{C}$ with center $\frac{1}{2}$ and radius $\frac{1}{2}$. Thus
$c_{\alpha}(1/2-i/2)\mathbf{1}-V_{n}-Z_{n}\notin G(A)\mbox{ and }c_{\alpha}(1/2-i/2)\mathbf{1}-V_{n}\in G(A),$
which together implies that
(3) $\mathbf{1}-Z_{n}\left(c_{\alpha}(1/2-i/2)\mathbf{1}-V_{n}\right)^{-1}\notin G(A).$
Since $V_{n}$ is normal we obtain the estimate
$s\left(\left(c_{\alpha}(1/2-i/2)\mathbf{1}-V_{n}\right)^{-1}\right)$ $=\rho\left(\left(c_{\alpha}(1/2-i/2)\mathbf{1}-V_{n}\right)^{-1}\right)$
$\leq\frac{1}{\operatorname{dist}(C_{r},\{c_{\alpha}(1/2-i/2)\})}$
from which it follows that $\lim_{n}s\left(Z_{n}\left(c_{\alpha}(1/2-i/2)\mathbf{1}-V_{n}\right)^{-1}\right)=0$, contradicting (3). Subsequently $c_{\alpha}\in\{0,1\}$, and $\phi(\alpha\mathbf{1}+x)\in\{0,\alpha\}$ follows as advertised. ∎
The proof of Theorem 3.8 now follows as in [11]:
Theorem 3.8.
Let $\phi$ be a continuous spectrally multiplicative function on $A$ with $\phi(\mathbf{1})=1$. Then $\phi(x)=\psi_{\phi}(x)$ for all $x$ in $A$, and hence $\phi$ is a character of $A$.
Proof.
For $x\in A$ define $K_{x}:=\{\alpha\in\mathbb{C}:\phi(\alpha\mathbf{1}+x)=0\}$, and assume first that $\psi_{\phi}(x)=0$. Our aim is to prove that $K_{x}=\{0\}$. Observe that $0\in K_{x}$ (by Lemma 3.5), $K_{x}\subseteq\sigma(-x)$, and, since $\phi$ is continuous, that $K_{x}$ is closed. Thus $K_{x}$ is nonempty and compact. Let $m$ be a maximum modulus element of $K_{x}$. From the definition of $m$ there is a sequence $(k_{n})\subset\mathbb{C}\setminus K_{x}$ which converges to $m$. Therefore, by Lemma 3.7, $\lim_{n}\phi\big{(}k_{n}\mathbf{1}+x\big{)}=\lim_{n}k_{n}=m$, and by continuity of $\phi$, $\lim_{n}\phi\big{(}k_{n}\mathbf{1}+x\big{)}=\phi\big{(}m\mathbf{1}+x\big{)}=0$. Thus $m=0$ from which it follows that $K_{x}=\{0\}$. Invoking Lemma 3.7 again we then obtain $\phi(\alpha\mathbf{1}+x)=\alpha$ for each $\alpha\in\mathbb{C}$. For any value of $\psi_{\phi}(x)$ we use the first part of the proof to deduce that
$\phi(x)=\phi\left(\psi_{\phi}(x)\mathbf{1}+\left[x-\psi_{\phi}(x)\mathbf{1}\right]\right)=\psi_{\phi}(x).$
As a direct consequence of Theorem 3.8 we then obtain:
Corollary 3.9.
Let $\phi$ be a continuous spectrally multiplicative function on $A$. Then either $\phi$ is a character of $A$ or $\phi$ is a character of $A$.
Proof.
If $\phi$ is spectrally multiplicative then so is $\phi$. But either $\phi\left(\mathbf{1}\right)=1$ or $\phi\left(\mathbf{1}\right)=1$. So the result follows from Theorem 3.8. ∎